A metal object at 54.0 oC is heated by gaining 30.2 kJ of heat from the...

  1. A metal object at 54.0 oC is heated by gaining 30.2 kJ of heat
    from the environment.
    The heat capacity, C, of the metal object is 6.30 kJ/ oC
  2. When 11.53 g C2H6O (g) is reacted according to the following equation, 368 kJ of heat is released.   What is the ΔH of the reaction?
    C2H6O (g)            +   3O2(g)        -->     2CO2(g)   +   3H2O (l)
  3. When 2.45 g of CH3OCH3 is burned at constant Pressure, 77.42 kJ of heat is given off. What is ∆H for this reaction? Hint: watch sign!
      
    CH3OCH3(l) + 3O2(g) -->    2CO3(s)    + 3H2O (g)    ∆ H rxn = ?
  4. SiO2(s) + 3C (graphite) --> SiC (s) + 2CO (g)      ΔH rxn =   626 kJ / mol rxn.
    Using the following standard enthalpy of formation data, calculate the standard enthalpy of formation for SiC (s).
    standard enthalpy of formation SiO2(s)   =     -910.9 kJ/mol
    standard enthalpy of formation CO (g) =  -110.5 kJ/mol
Answer #1

1)

Q = Cp dT

30.2 = 6.30 x (Tf - 54)

Tf = 58.8

final temperature = 58.8 oC

2)


C2H6O (g)            +   3O2(g)        -->     2CO2(g)   +   3H2O (l)

moles of C2H6O = 11.53 / 46 = 0.250 mol

ΔH = Q / n

     = 368 / 0.250

ΔH = -1468 kJ/mol

3)
  
CH3OCH3(l) + 3O2(g) -->    2CO3(s)    + 3H2O (g)    ∆ H rxn = ?

moles of CH3OCH3 = 2.45 / 46 = 0.0533 mol

∆H = - 77.42 / 0.0533

∆ H rxn = - 1453 kJ / mol

4)

standard enthalpy of formation for SiC (s). = - 63.9 kJ/mol